OpenStudy (anonymous):

What is the distance between the coordinates (5, 5) and (7, 2)?

OpenStudy (anonymous):

\[d = \sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\]

OpenStudy (anonymous):

distance formula is \[d = \sqrt{(x _{2} - x _{1})^2 + (y _{2} - y _{1})^2}\]

OpenStudy (anonymous):

dammit!

OpenStudy (anonymous):

\[d = \sqrt{(7 - 5)^{2} + (2 - 5)^{2}}\]lol. Gotta work faster :)

OpenStudy (anonymous):

I know that, i've done the quiz 3 times im on my 4th try!! But every time i do the formula, then i get all confused....and i mess up even when im not!! D:

OpenStudy (anonymous):

\[d = \sqrt{(2)^{2} + (-3)^{2}}\]\[d = \sqrt{4 + 9}\]\[d = \sqrt{13}\] from there, you can plug it into a calculator for a approximate answer or leave it like that for an exact answer.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yes thank you! <3